# Classification of compact manifolds with positive isotropic curvature

### v6

@@ Abstract
We show the following result: Let $(M,g_{0})$ be a compact connected manifold of dimension $n\geq 12$ with positive isotropic curvature. Then $M$ is diffeomorphic to a spherical space form, or a quotient manifold of $\mathbb{S}^{n-1}\times\mathbb{R}$ by a cocompact discrete subgroup of the isometry group of the round cylinder $\mathbb{S}^{n-1}\times\mathbb{R}$, or a connected sum of a finite number of such manifolds. This extends previous works of Brendle and Chen-Tang-Zhu, and improves a work of Huang. The proof uses Ricci flow with surgery on compact orbifolds, with the help of the ambient isotopy uniqueness of closed tubular neighborhoods of compact suborbifolds.
@@ 1 Introduction
(Of course, the cocompact discrete subgroup of the isometry group of the round cylinder $\mathbb{S}^{n-1}\times\mathbb{R}$ referred to in the above theorem should act freely on $\mathbb{S}^{n-1}\times\mathbb{R}$; moreover, it is well known that $\text{Isom}(\mathbb{S}^{n-1}\times\mathbb{R})=\text{Isom}(\mathbb{S}^{n-1})\times\text{Isom}(\mathbb{R})$ for $n\geq 3$. By [24] the converse of Theorem 1.1 is also true.) This extends the main theorem in Chen-Tang-Zhu [10], which deals with the 4 dimensional case, and Theorem 1.4 in Brendle [7], which deals with the case that the $n$-manifolds contain no nontrivial incompressible $(n-1)$-dimensional space forms, where $n\geq 12$. The result is analogous to the classification of compact 3-manifolds with positive scalar curvature by Perelman et al.
@@
The following result is a simple consequence of Theorem 1.1; compare the Main Theorem in [23].
@@
In particular, since any compact manifold which admits a non-collapsed, ancient Ricci flow must have finite fundamental group, cf. [1], it follows that any compact (connected) manifold of dimension $n\geq 12$ which admits a non-collapsed, ancient Ricci flow with positive isotropic curvature is diffeomorphic to a spherical space form. (Recall that by combining Perelman’s result that “there is no shrinking breathers other than gradient solitons” on closed manifolds (see p.9 of [32], cf. also Theorem 1.3 in [30]) with Theorem 1.4 in [31] one sees that any compact shrinking Ricci soliton of dimension $n\geq 12$ with positive isotropic curvature is isometric to a spherical space form.)
@@
Corollary 1.5.
Let $M$ be a compact smooth manifold of dimension $n\geq 12$ which admits a metric with positive isotropic curvature. Then $\pi_{i}(M)=0$ for $22\leq i\leq n-$ 2.
In [24] Micallef-Wang showed that a compact manifold $M$ of even dimension $n=2m\geq 4$ with positive isotropic curvature has the second Betti number $b_{2}(M)=0$. The following corollary extends their result when $n\geq 12$ (for our result the dimension $n$ need not be even).
Corollary 1.6.
Let $M$ be a compact smooth manifold of dimension $n\geq 12$ which admits a metric with positive isotropic curvature. Then $b_{i}(M)=0$ for $22\leq i\leq n-$ 2.
@@
(By an obvious extension of Theorem 1.1 in [24] to the case of orbifold connected sum (cf. also [9]), certain version of the converse of Theorem 1.7 is also true.) This improves Theorem 1.3 in [19]. Compare Theorem 2.1 in [10]. We emphasize that here our notion of orbifold connected sums (see [19]) is stronger than that defined on p. 48 of [10]. (As in [10] we also allow self-connected sums; but of course our notion of self-connected sums is also stronger than that given in [10].)
As in [10], [7], and [19], the main tool in our proof is Ricci flow with surgery, aided by some topological arguments. In the 4-dimensional case the topological part of the proof of the main theorem in Chen-Tang-Zhu [10] used a result of McCullough [26], which says that any diffeomorphism of a 3-dimensional spherical space form is isotopic to an isometry. Similar statement for $(n-1)$-dimensional spherical space forms with $n\geq 5$ is not true in general.
One of our key observations is that for the topological part in the higher dimensional case, it suffices to use the ambient isotopy uniqueness of closed tubular neighborhoods of compact suborbifolds. More precisely, using this uniqueness and a (rough) classification of spherical orbifolds with (nonempty) isolated singularities (due to [10] in the even ($\geq 4$) dimensional case and to [19] in the odd ($\geq 3$) dimensional case) one can prove fine properties of orbifold ancient $\kappa$-solutions and standard solutions when they contain caps with ends of type $\mathbb{S}^{n-1}/\Gamma\times(-1,1)$ with $\Gamma$ nontrivial, extending what were done in [7] and [19] for caps with ends of type $\mathbb{S}^{n-1}\times(-1,1)$, and extend/strengthen the definition of $\varepsilon$-caps in [7] and [19]. Then we can recognize the topology of components in the surgery process which are covered by canonical neighborhoods. Moreover, using the ambient isotopy uniqueness of closed tubular neighborhoods of compact suborbifolds again, one can determine the topology of a connected sum of at most two spherical orbifolds with at most isolated singularities. (It is clear from our arguments in this paper that the use of McCullough’s result in the proof of the main theorem in Chen-Tang-Zhu [10] can be avoided; moreover, to get a similar classification of compact manifolds of dimension $5\leq n\leq 11$ with positive isotropic curvature, the only thing that remains to do is to show a suitable curvature pinching estimate, as Hamilton did in dimension 4 in [16] and Brendle did in dimensions $n\geq 12$ in [7].)
@@
In Section 2 we define $\varepsilon$-caps, and describe the canonical neighborhood structure of orbifold ancient $\kappa$-solutions and orbifold standard solutions. In Section 3 we construct an $(r,\delta)$-surgical solution to the Ricci flow starting with a compact, connected Riemannian orbifold of dimension $n\geq 12$ with at most isolated singularities and with positive isotropic curvature. As applications we prove Theorems 1.7 and 1.1, Corollaries 1.2, 1.3, 1.5 and 1.6, and derive a topological classification of compact Riemannian orbifold of dimension $n\geq 12$ with at most isolated singularities and with positive isotropic curvature. We will follow the notation and conventions in [19] in general.
@@ 2 $\varepsilon$-caps and ancient $\kappa$-solutions
To describe the structure of orbifold ancient $\kappa$-solutions and orbifold standard solutions we need notions on necks and caps.
Suppose $\sigma$ is an isometric involution of the spherical space form $\mathbb{S}^{n-1}/\Gamma$ with at most isolated fixed points, let $\hat{\sigma}$ be the involution on the manifold $\mathbb{S}^{n-1}/\Gamma\times\mathbb{R}$ defined by $\hat{\sigma}(x,s)=(\sigma(x),-s)$ for $x\in\mathbb{S}^{n-1}/\Gamma$ and $s\in\mathbb{R}$, consider the quotient orbifold $(\mathbb{S}^{n-1}/\Gamma\times\mathbb{R})//\langle\hat{\sigma}\rangle$, which has at most isolated singularities. We also denote this (Riemannian) orbifold by $\mathbb{S}^{n-1}/\Gamma\times_{\mathbb{Z}_{2}}\mathbb{R}$; but beware that this notation has some ambiguity. Similarly we can define $\mathbb{S}^{n-1}/\Gamma\times_{\mathbb{Z}_{2}}[-1,1]$. Note that we can consider $\Gamma$ and $\hat{\sigma}$ as isometries of $\mathbb{S}^{n}$ in a natural way, by lifting $\sigma$ to an isometry of $\mathbb{S}^{n-1}$ (which is always possible) and viewing $\mathbb{S}^{n}$ as a suspension of $\mathbb{S}^{n-1}$. We’ll use the same notation for these isometries of $\mathbb{S}^{n}$. The $(\mathbb{S}^{n-1}/\Gamma\times\mathbb{R})//\langle\hat{\sigma}\rangle$ above is a smooth manifold if and only if $\sigma$ has no any fixed points in $\mathbb{S}^{n-1}/\Gamma$; if this is the case, it is diffeomorphic to $\mathbb{S}^{n}//\langle\Gamma,\hat{\sigma}\rangle\setminus\bar{B}$, where $B$ is a small, open metric ball centered at the point coming from the north and south poles of $\mathbb{S}^{n}$. Note that when $n$ is odd, $(\mathbb{S}^{n-1}/\Gamma\times\mathbb{R})//\langle\hat{\sigma}\rangle$ is a smooth manifold if and only if $\Gamma$ is trivial and $\langle\sigma\rangle$ acts on $\mathbb{S}^{n-1}$ antipodally; see the proof of Proposition 2.5 in [19]. Below when we write $\mathbb{S}^{n}//\langle\Gamma,\hat{\sigma}\rangle$, we always assume that the involution $\sigma$ has no any fixed points in $\mathbb{S}^{n-1}/\Gamma$.
@@
Let $\mathcal{O}$ be a $n$-dimensional orbifold, $U$ an open subset, and $\Gamma$ as above. If $U$ is diffeomorphic to $\mathbb{S}^{n-1}//\Gamma\times\mathbb{R}$, we call it a topological neck. If $U$ is diffeomorphic to $\mathbb{R}^{n}//\Gamma$ or $\mathbb{S}^{n-1}//\Gamma\times_{\mathbb{Z}_{2}}\mathbb{R}$, we call it a topological cap.
@@
Let $\mathcal{O}$ be an orbifold, and $\mathcal{O}^{\prime}$ be a suborbifold of $\mathcal{O}$ in the sense of p. 108 in [21], so there is an immersion $f:\mathcal{O}^{\prime}\rightarrow\mathcal{O}$ such that $|f|$ maps $|\mathcal{O}^{\prime}|$ homeomorphically onto its image in $|\mathcal{O}|$. We identify $\mathcal{O}^{\prime}$ with its image in $\mathcal{O}$. Let $E(\xi)$ be the total space of an orbifold vector bundle $\xi$ over $\mathcal{O}^{\prime}$, and $\varphi:E(\xi)\rightarrow\mathcal{O}$ be a diffeomorphism onto an open neighborhood of $\mathcal{O}^{\prime}$ in $\mathcal{O}$, with the restriction of $\varphi$ to the zero-section of $\xi$ coinciding with the inclusion of $\mathcal{O}^{\prime}$ in $\mathcal{O}$. We call $\varphi$ an open tubular neighborhood of $\mathcal{O}^{\prime}$. Now we endow $\xi$ with a fiberwise inner product structure, so that it becomes a Euclidean orbifold vector bundle. Then we call the restriction of the above diffeomorphism $\varphi$ to the (unit) disk bundle $D(\xi)$ of $\xi$ a closed tubular neighborhood of $\mathcal{O}^{\prime}$. Compare Section 2 in Chapter VI of [6], Section 2 in Chapter III of [22], and Chapter 7 in [29].
@@
Two open tubular neighborhoods $\varphi:E(\xi)\rightarrow\mathcal{O}$ and $\psi:E(\eta)\rightarrow\mathcal{O}$ of $\mathcal{O}^{\prime}$ is isotopic if there are a family of open tubular neighborhoods $\varphi_{t}:E(\xi)\rightarrow\mathcal{O}$ of $\mathcal{O}^{\prime}$ and a vector bundle isomorphism $\Lambda:E(\xi)\rightarrow E(\eta)$ such that $\varphi_{0}=\varphi$, $\varphi_{1}=\psi\circ\Lambda$, and the map $[0,1]\times E(\xi)\rightarrow\mathcal{O}$ taking $(t,v)\mapsto\varphi_{t}(v)$ is smooth. If $\xi$ and $\eta$ are endowed with fiberwise inner product structures, and the above $\Lambda$ can be taken to be an (orthogonal) isomorphism of Euclidean orbifold vector bundles, then we call the corresponding closed tubular neighborhoods of $\mathcal{O}^{\prime}$ isotopic. Compare Section 2 in Chapter VI of [6] and Chapter 7 in [29].
@@
The following result on ambient isotopy uniqueness of closed tubular neighborhoods of a compact suborbifold in an orbifold should be known; cf. p.443 of [5]. It will be used extensively in this paper.

### v8

@@ Abstract
We show the following result: Let $(M,g_{0})$ be a compact connected manifold of dimension $n\geq 12$ with positive isotropic curvature. Then $M$ is diffeomorphic to a spherical space form, or a quotient manifold of $\mathbb{S}^{n-1}\times\mathbb{R}$ by a cocompact discrete subgroup of the isometry group of the round cylinder $\mathbb{S}^{n-1}\times\mathbb{R}$, or a connected sum of a finite number of such manifolds. This extends previous works of Brendle and Chen-Tang-Zhu, and improves a work of Huang. The proof uses Ricci flow with surgery on compact orbifolds, with the help of the ambient isotopy uniqueness of closed tubular neighborhoods of compact embedded full suborbifolds.
@@ 1 Introduction
(For the classification of spherical space forms see [51]. Of course, the cocompact discrete subgroup of the isometry group of the round cylinder $\mathbb{S}^{n-1}\times\mathbb{R}$ referred to in the above theorem should act freely on $\mathbb{S}^{n-1}\times\mathbb{R}$; moreover, it is well known that $\text{Isom}(\mathbb{S}^{n-1}\times\mathbb{R})\cong\text{Isom}(\mathbb{S}^{n-1})\times\text{Isom}(\mathbb{R})$ for $n\geq 2$ (cf. for example, the proof of Corollary 6.2 in [13]). For the investigation of compact quotient manifolds of $\mathbb{S}^{n-1}\times\mathbb{R}$ by standard isometries one can consult for example, p. 457 of [48], p. 827 of [14], Theorem 1.1 in [23], and [32]. By [34] the converse of Theorem 1.1 is also true.) This extends the Main Theorem in Chen-Tang-Zhu [15], which deals with the 4 dimensional case, and Theorem 1.4 in Brendle [11], which deals with the case that the $n$-manifolds contain no nontrivial incompressible (n1)(n-1)-dimensional space forms, where $n\geq 12$. The result is analogous to the classification of compact 3-manifolds with positive scalar curvature by Perelman et al.
@@
The following result is a simple consequence of Theorem 1.1, which strengthens and extends the Main Theorem in Micallef-Moore [33] in the case of $n\geq 12$.
@@
In particular, since any compact manifold which admits a non-collapsed, ancient Ricci flow must have finite fundamental group, cf. [2], it follows that any compact (connected) manifold of dimension $n\geq 12$ which admits a non-collapsed, ancient Ricci flow with positive isotropic curvature is diffeomorphic to a spherical space form. (Recall that by combining Perelman’s result that “there is no shrinking breathers other than gradient solitons” on a closed manifold (see p.9 of [42], cf. also Theorem 1.3 in [40]) with Theorem 1.4 in [41] one sees that any compact shrinking Ricci soliton of dimension $n\geq 12$ with positive isotropic curvature is isometric to a spherical space form.)
Corollary 1.4.
Let $M$ be a compact smooth manifold of dimension $n\geq 12$ which admits a metric with positive isotropic curvature. Then $\pi_{i}(M)=0$ for $2\leq i\leq n-2$.
In [34] Micallef-Wang showed that a compact manifold $M$ of even dimension $n=2m\geq 4$ with positive isotropic curvature has the second Betti number $b_{2}(M)=0$. The following corollary extends their result when $n\geq 12$ (for our result the dimension $n$ needs not be even).
Corollary 1.5.
Let $M$ be a compact smooth manifold of dimension $n\geq 12$ which admits a metric with positive isotropic curvature. Then $b_{i}(M)=0$ for $2\leq i\leq n-2$.
@@
(By an obvious extension of Theorem 1.1 in [34] to the case of orbifold connected sum (cf. also [14]), a certain version of the converse of Theorem 1.7 is also true.) This improves Theorem 1.3 in [27]. Compare Theorem 2.1 in [15]. We emphasize that here our notion of orbifold connected sums (see [27]) is stronger than that defined on p. 48 of [15]. (As in [15] we also allow self-connected sums; but of course our notion of self-connected sums is also stronger than that given in [15].)
As in [15], [11], and [27], the main tool in our proof is Ricci flow with surgery, aided by some topological arguments. In the 4-dimensional case the topological part of the proof of the Main Theorem in Chen-Tang-Zhu [15] used a result of McCullough [36], which says that any diffeomorphism of a 3-dimensional spherical space form is isotopic to an isometry. Similar statement for (n1)(n-1)-dimensional spherical space forms with $n\geq 5$ is not true in general.
One of our key observations is that for the topological part in the higher dimensional case, it suffices to use the ambient isotopy uniqueness of closed tubular neighborhoods of compact embedded full suborbifolds (for the definitions see Section 2). More precisely, using this uniqueness and a (rough) classification of spherical orbifolds with (nonempty) isolated singularities (due to [15] in the even ($\geq 4$) dimensional case and to [27] in the odd ($\geq 3$) dimensional case) one can prove fine properties of orbifold ancient $\kappa$-solutions and standard solutions when they contain caps with ends of type $\mathbb{S}^{n-1}/\Gamma\times(-1,1)$ with $\Gamma$ nontrivial, extending what were done in [11] and [27] for caps with ends of type $\mathbb{S}^{n-1}\times(-1,1)$, and extend/strengthen the definition of $\varepsilon$-caps in [11] and [27]. Then we can recognize the topology of components in the surgery process which are covered by canonical neighborhoods. Moreover, using the ambient isotopy uniqueness of closed tubular neighborhoods of compact embedded full suborbifolds again, one can determine the topology of a connected sum of at most two spherical orbifolds with at most isolated singularities. (It is clear from our arguments in this paper that the use of McCullough’s result in the proof of the Main Theorem in Chen-Tang-Zhu [15] can be avoided; moreover, to get a similar classification of compact manifolds of dimension $5\leq n\leq 11$ with positive isotropic curvature, the only thing that remains to do is to show a suitable curvature pinching estimate, as Hamilton did in dimension 4 in [24] and Brendle did in dimensions $n\geq 12$ in [11].)
In fact, in this paper we only use the tubular neighborhood theorem (Proposition 2.2) in the case that the ambient orbifold is a spherical orbifold with at most isolated singularities. In this case the tubular neighborhood theorem is implied by an equivariant tubular neighborhood theorem for an ambient manifold diffeomorphic to either $\mathbb{S}^{n}$ or $\mathbb{R}^{n}$ and with a finite group action, which is well established in the existing literature. But we’ll show a more general orbifold version of the tubular neighborhood theorem, whose precise statement and detailed proof are not easy to find in the previous literature. We find that it is very convenient to use possibly noneffective orbifolds and orbifold vector bundles over possibly noneffective orbifolds here.
@@
In Section 2 we show the ambient isotopy uniqueness of closed tubular neighborhoods of compact embedded full suborbifolds, define $\varepsilon$-caps, and describe the canonical neighborhood structure of orbifold ancient $\kappa$-solutions and orbifold standard solutions. In Section 3 we construct an $(r,\delta)$-surgical solution to the Ricci flow starting with a compact, connected Riemannian orbifold of dimension $n\geq 12$ with at most isolated singularities and with positive isotropic curvature. As applications we prove Theorems 1.7 and 1.1, Corollaries 1.2, 1.3, 1.4 and 1.5, and derive a topological classification of compact Riemannian orbifold of dimension $n\geq 12$ with at most isolated singularities and with positive isotropic curvature. We will follow the notation and conventions in [27] in general. In particular, as in [27], when necessary we’ll assume that the smooth maps between orbifolds that appear are smooth complete orbifold maps in the sense of [7] and [9] to avoid any possible ambiguity.
@@ 2 $\varepsilon$-caps and ancient $\kappa$-solutions
To describe the structure of orbifold ancient $\kappa$-solutions and orbifold standard solutions we need notions on necks and caps. Let $\mathcal{O}$ be an $n$-dimensional orbifold, $U$ an open subset in $\mathcal{O}$ (sometimes when we write $\mathcal{O}$ we really mean its underlying space $|\mathcal{O}|$). If $U$ is diffeomorphic to $\mathbb{S}^{n-1}//\Gamma\times\mathbb{R}$, we call it a topological neck.
Suppose further that the spherical space form $\mathbb{S}^{n-1}/\Gamma$ admits an isometric involution $\sigma$ with at most isolated fixed points, let $\hat{\sigma}$ be the involution on the manifold $\mathbb{S}^{n-1}/\Gamma\times\mathbb{R}$ defined by $\hat{\sigma}(x,s)=(\sigma(x),-s)$ for $x\in\mathbb{S}^{n-1}/\Gamma$ and $s\in\mathbb{R}$, consider the quotient orbifold $(\mathbb{S}^{n-1}/\Gamma\times\mathbb{R})//\langle\hat{\sigma}\rangle$, which has at most isolated singularities. We also denote this (Riemannian) orbifold by $\mathbb{S}^{n-1}/\Gamma\times_{\mathbb{Z}_{2}}\mathbb{R}$; but beware that this notation has some ambiguity. Similarly we can define $\mathbb{S}^{n-1}/\Gamma\times_{\mathbb{Z}_{2}}[-1,1]$. Note that we can consider $\Gamma$ and $\hat{\sigma}$ as isometries of $\mathbb{S}^{n}$ in a natural way, by lifting $\sigma$ to an isometry of $\mathbb{S}^{n-1}$ (which is always possible) and viewing $\mathbb{S}^{n}$ as a suspension of $\mathbb{S}^{n-1}$. We’ll use the same notation for these isometries of $\mathbb{S}^{n}$. The $(\mathbb{S}^{n-1}/\Gamma\times\mathbb{R})//\langle\hat{\sigma}\rangle$ above is a smooth manifold if and only if $\sigma$ has no any fixed points in $\mathbb{S}^{n-1}/\Gamma$; if this is the case, it is diffeomorphic to $\mathbb{S}^{n}//\langle\Gamma,\hat{\sigma}\rangle\setminus\bar{B}$, where $B$ is a small, open metric ball centered at the point coming from the north and south poles of $\mathbb{S}^{n}$. Note that when $n$ is odd, $(\mathbb{S}^{n-1}/\Gamma\times\mathbb{R})//\langle\hat{\sigma}\rangle$ is a smooth manifold if and only if $\Gamma$ is trivial and $\langle\sigma\rangle$ acts on $\mathbb{S}^{n-1}$ antipodally; see the proof of Proposition 2.5 in [27]. Below when we write $\mathbb{S}^{n}//\langle\Gamma,\hat{\sigma}\rangle$, we always assume that the involution $\sigma$ has no any fixed points in $\mathbb{S}^{n-1}/\Gamma$. (For the investigation of spherical space forms which admit free isometric involutions one can consult for example, [23] and [32].)
@@
If an open subset $U$ in an $n$-dimensional orbifold $\mathcal{O}$ is diffeomorphic to $\mathbb{R}^{n}//\Gamma$ (here $\Gamma$ is a finite subgroup of $O(n)$ acting freely on $\mathbb{S}^{n-1}$) or $\mathbb{S}^{n-1}//\Gamma\times_{\mathbb{Z}_{2}}\mathbb{R}$ (in this case the spherical space form $\mathbb{S}^{n-1}/\Gamma$ should admit an isometric involution $\sigma$ with at most isolated fixed points) defined above, we call it a topological cap.
Let $\mathcal{O}$ and $\mathcal{O}^{\prime}$ be two orbifolds, and $f:\mathcal{O}^{\prime}\rightarrow\mathcal{O}$ be an immersion (as defined on p. 107-108 of [29]) such that $|f|$ maps $|\mathcal{O}^{\prime}|$ homeomorphically onto its image in $|\mathcal{O}|$. We call $f$ or $\mathcal{O}^{\prime}$ an embedded suborbifold of $\mathcal{O}$ (but beware that the meaning of the word “embedded” here is not the same as that on p. 108 of [29]). We identify $\mathcal{O}^{\prime}$ with its image in $\mathcal{O}$. We assume further that the image of $\mathcal{O}^{\prime}$ under $f$ is a full suborbifold of $\mathcal{O}$ in the sense of Definition 5 in [9] (see also [50]; compare Definition 1.14 in [21]). We’ll call such an immersion $f$ (or embedded suborbifold $\mathcal{O}^{\prime}$) an embedded full suborbifold (see also [50] and Definition 2.5 in [31]); there is also an obvious extension of the notion of embedded full suborbifold to the case of embedded full suborbifold with boundary, in which the orbifold $\mathcal{O}^{\prime}$ above can have boundary, correspondingly the suborbifold $\mathcal{P}$ and the submanifold $\widetilde{V}_{x}$ in Definition 4 in [9] can have boundary. (Note that any point in $|\mathcal{O}|$ may be viewed as a 0-dimensional embedded full suborbifold of $\mathcal{O}$.)
Now we view the embedded full suborbifold $\mathcal{O}^{\prime}$ as a possibly noneffective orbifold with local groups (whose actions on local models may be noneffective) at all $p\in|\mathcal{O}^{\prime}|\subset|\mathcal{O}|$ the same as the ones in the ambient orbifold $\mathcal{O}$. We still denote this possibly noneffective orbifold by $\mathcal{O}^{\prime}$. Let $\xi$ be an orbifold vector bundle over the possibly noneffective orbifold $\mathcal{O}^{\prime}$ such that there is an embedding $\varphi:E(\xi)\rightarrow\mathcal{O}$ with $|\varphi|(|E(\xi)|)$ an open neighborhood of $|\mathcal{O}^{\prime}|$ in $|\mathcal{O}|$, and with the restriction of $\varphi$ to the zero-section of $\xi$ coinciding with the inclusion of $\mathcal{O}^{\prime}$ in $\mathcal{O}$, where $E(\xi)$ is the total space of the bundle $\xi$, and is required to be an effective orbifold. (The definition of an orbifold vector bundle over a possibly noneffective orbifold can be easily adapted from that of an orbivector bundle on p. 107 of [29]; the only difference is that for any point $p$ in the underlying space of the base orbifold, the local group $G_{p}$ may act noneffectivily on the local model of the base orbifold around $p$.) We call $\varphi$ an open tubular neighborhood of $\mathcal{O}^{\prime}$ in $\mathcal{O}$. For an embedded full suborbifold $\mathcal{O}^{\prime}$ (without boundary) of $\mathcal{O}$ as above, such an orbifold vector bundle $\xi$ always exists; for example one can choose $\xi$ to be the normal bundle of the suborbifold $\mathcal{O}^{\prime}$ in $\mathcal{O}$. Here we define the normal bundle of the embedded full suborbifold $\mathcal{O}^{\prime}$ in $\mathcal{O}$ to be the orbifold vector bundle $i^{*}T\mathcal{O}/T\mathcal{O}^{\prime}$ over the possibly noneffective orbifold $\mathcal{O}^{\prime}$, where $i:\mathcal{O}^{\prime}\rightarrow\mathcal{O}$ is the inclusion map (from the possibly noneffective orbifold $\mathcal{O}^{\prime}$ to the ambient orbifold $\mathcal{O}$) with all the corresponding homomorphisms between local groups being the identity maps; consult p. 35-37 of [1] and p. 304 of [21]. (The definition of the pullback of the tangent bundle used here can be obviously adapted from Definition 6.2 in [8]; the only difference is that for any point $x$ in the underlying space of the domain orbifold, the local group $\Gamma_{x}$ (using the same notation as in [8]) may act noneffectivily on the local model of the domain orbifold around $x$.) One can show this by adapting the proof of Theorem 12.11 in [12] and Theorem 7.1.5 in [39] to the orbifold case; consult also Theorem 2.2 in Chapter VI of [10], where the closeness condition imposed on the invariant submanifold $A$ can be removed, Proposition 2.1.2 in [20] and Proposition 1.15 in [21], whose proof uses Theorem 2.2 in Chapter VI of [10] and the orthonormal frame bundle of $\mathcal{O}$ (after introducing a Riemannian metric on $\mathcal{O}$), Theorem 7.3 in [28], and Theorem 2.10 in [44].
Example. Let $\mathcal{O}^{\prime}$ be a point (a 0-dimensional connected manifold), $\mathcal{O}$ be an orbifold of dimension $n\geq 1$. Suppose $q\in|\mathcal{O}|$ is a singular point with local group $G_{q}$, and $f:\mathcal{O}^{\prime}\rightarrow\mathcal{O}$ is the map with image $\{q\}$. Then $f$ (or $\mathcal{O}^{\prime}$ identified with $\{q\}$) is an embedded full suborbifold of $\mathcal{O}$. By the above definition the normal bundle of $\mathcal{O}^{\prime}$ in $\mathcal{O}$ is isomorphic to $\mathbb{R}^{n}//G_{q}$, which is an orbifold vector bundle over a 0-dimensional (connected) noneffective orbifold with local group $G_{q}$.
Example (see Example 10 in [9]). Let $\mathcal{Q}$ be the quotient orbifold $\mathbb{R}//\mathbb{Z}_{2}$, where the $\mathbb{Z}_{2}$-action on $\mathbb{R}$ is generated by the reflection with respect to the origin. Let $\Delta:\mathcal{Q}\rightarrow\mathcal{Q}\times\mathcal{Q}$ be the diagonal embedding. Note that $\Delta$ is not a full suborbifold of $\mathcal{Q}\times\mathcal{Q}$. If we still define the normal bundle of the immersion $\Delta$ to be $N_{\Delta}:=\Delta^{*}T(\mathcal{Q}\times\mathcal{Q})/T\mathcal{Q}$ (but here we view $\mathcal{Q}$ as an effective orbifold), then $N_{\Delta}$ is isomorphic to the tangent bundle $T\mathcal{Q}$, and the only singular point of the total space of the bundle $N_{\Delta}$ has local group $\mathbb{Z}_{2}$. But any neighborhood of the image of $\Delta$ (we also denote this image by $\Delta$) in $\mathcal{Q}\times\mathcal{Q}$ has a singular point with local group $\mathbb{Z}_{2}\times\mathbb{Z}_{2}$. One sees that the suborbifold $\Delta$ admits no reasonable tubular neighborhood in $\mathcal{Q}\times\mathcal{Q}$.
@@
Now we endow the bundle $\xi$ above with a fiberwise inner product structure, so that it becomes a Euclidean orbifold vector bundle. Then we call the restriction of the above diffeomorphism $\varphi$ to the (unit) disk bundle $D(\xi)$ of $\xi$ a closed tubular neighborhood of $\mathcal{O}^{\prime}$. Compare Section 2 in Chapter VI of [10], Section 2 in Chapter III of [30], and Chapter 7 in [39].
@@
Two open tubular neighborhoods $\varphi:E(\xi)\rightarrow\mathcal{O}$ and $\psi:E(\eta)\rightarrow\mathcal{O}$ of $\mathcal{O}^{\prime}$ are isotopic if there exist a family of open tubular neighborhoods $\varphi_{t}:E(\xi)\rightarrow\mathcal{O}$ of $\mathcal{O}^{\prime}$ and a vector bundle isomorphism $\Lambda:E(\xi)\rightarrow E(\eta)$ such that $\varphi_{0}=\varphi$, $\varphi_{1}=\psi\circ\Lambda$, and the map $0,1]\times E(\xi)\rightarrow\mathcal{O}$ taking $(t,v)\mapsto\varphi_{t}(v)$ is smooth. If $\xi$ and $\eta$ are endowed with fiberwise inner product structures, and the above $\Lambda$ can be taken to be an (orthogonal) isomorphism of Euclidean orbifold vector bundles, then we call the corresponding closed tubular neighborhoods of $\mathcal{O}^{\prime}$ isotopic. Compare Section 2 in Chapter VI of [10] and Chapter 7 in [39].
We need a simple version of Thom’s isotopy extension theorem for orbifolds.
Lemma 2.1.
Let $\mathcal{P}$ be a compact orbifold with or without boundary, $\mathcal{O}$ be an orbifold (without boundary), and $f:\mathcal{P}\rightarrow\mathcal{O}$ be an embedded full suborbifold. Suppose that $h:\mathcal{P}\times[0,1]\rightarrow\mathcal{O}$ is an isotopy of $f$ in the sense of Definition 9.1 in [12] (extended to the orbifold case) such that each $h_{t}:=h(\cdot,t)$ is an embedded full suborbifold. Then there exists a diffeotopy $H:\mathcal{O}\times[0,1]\rightarrow\mathcal{O}$ of $\mathcal{O}$ in the sense of Definition 9.3 in [12] (extended to the orbifold case) with compact support (in the sense of p. 207 of [39]) such that $h_{t}=H_{t}\circ h_{0}$ for all $t\in[0,1]$.
Proof. The proof can be easily adapted from that of Theorem 9.5 in [12] and Theorem 7.3.3 in [39] (cf. also p. 179-181 of [35] and the proof of Theorem 8.6 in [28] and Theorem 2.4.2 in [49]). More precisely, the desired ambient isotopy comes from the projection of a suitable global flow on $\mathbb{R}\times\mathcal{O}$, the latter being generated by a vector field on $\mathbb{R}\times\mathcal{O}$. To construct this vector field, we first construct suitable vector fields in local models of $\mathbb{R}\times\mathcal{O}$ as in the proof of Theorem 9.5 in [12], Theorem 7.3.3 in [39] and Theorem 2.4.2 in [49], then get invariant vector fields in local models by using the method of averaging as in the proof of Theorem 3.1 in Chapter VI in [10], finally glue these local vector fields to a global vector field via a partition of unity. $\Box$
@@
The following result on ambient isotopy uniqueness of closed tubular neighborhoods of a compact embedded full suborbifold in an orbifold should be known; cf. p.443 of [6]. (Compare also the Remark on p.312 of [10] and Theorem 2.9 in [44].) It will be used extensively in this paper.